### v1

# Early-Stage Anomaly Detection: A Study of Model Performance on Complete vs. Partial Flows

Abstract
This study investigates the efficacy of machine learning models, specifically Random Forest, in anomaly detection systems when trained on complete flow records and tested on partial flow data. We explore the performance disparity that arises when models are applied to incomplete data typical in real-world, real-time network environments. Our findings demonstrate a significant decline in model performance, with precision and recall dropping by up to 30% under certain conditions when models trained on complete flows are tested against partial flows. Conversely, models trained and tested on consistently complete or partial datasets maintain robustness, highlighting the importance of dataset consistency in training. The study reveals that a minimum of 7 packets in the test set is required for maintaining reliable detection rates. These results underscore the need for tailored training strategies that can effectively adapt to the dynamics of partial data, enhancing the practical applicability of anomaly detection systems in operational settings.
Index Terms:
early-stage anomaly detection, cybersecurity, complete flow information, partial flow information.
I Introduction
In the rapidly evolving field of network security, the effectiveness of machine learning (ML)-based anomaly detection systems is increasingly being tested in real-time environments. A fundamental limitation in current research is the gap between the conditions under which ML models are developed and validated, and the dynamic nature of real-world network operations. This paper addresses a critical aspect often overlooked: the disparity between models trained on complete flow records and the practical necessity of real-time anomaly detection using partial or evolving flow information.
Existing research predominantly relies on datasets generated to capture comprehensive flow records, encompassing the full lifecycle of network flows [7307098, 8543584, Kadri2024]. This approach, while detailed, does not accurately reflect the transient and incomplete nature of data in real-time network monitoring. In real-world scenarios, anomaly detection systems must often make swift decisions based on incomplete information—a stark contrast to the complete flow records used in standard dataset configurations. This discrepancy raises significant concerns about the transferability of research findings to real-time applications, where the speed and accuracy of anomaly detection are paramount.
To address this disconnect, our work evaluates how the performance of machine learning (ML) models, specifically Random Forest (RF), varies under different training and testing conditions. We experiment with models trained on complete flow records and tested on incomplete ones, as well as models trained on incomplete data and tested on the same. To simulate these scenarios, we define specific thresholds on packet counts and flow durations, assessing the effectiveness of the models with both fully and partially captured datasets.
Random Forest is selected for this investigation due to its common use in cybersecurity and anomaly detection, as demonstrated in [CICIDS2017, engelen2021, lanvin2023]. By incorporating RF into our analysis, we establish a common baseline for comparability, facilitating a more grounded evaluation of the models’ effectiveness in handling incomplete data sets in security contexts.
Our findings reveal that models trained on complete flows and tested on partial flows experience a notable decline in performance, with precision and recall dropping by up to 30% under certain conditions. Conversely, except for some isolated cases, models trained and tested consistently on either complete or partial flows maintain robustness, illustrating the critical impact of dataset consistency on model efficacy. This disparity emphasizes the challenge of transferring lab-based model accuracy to real-world, real-time detection systems, where data incompleteness can markedly impact detection capabilities. Our study also demonstrates that, for the dataset examined, at least 7 packets in the test set are required to maintain acceptable detection rates in real-time scenarios, underlining the need for model adjustments or enriched training strategies to handle partial data effectively.
The rest of this paper is organized as follows: Section II discusses the selection of thresholds for model evaluation, including an overview of the used dataset and the ML algorithm employed. Section III details the process of raw data preprocessing, flow measurement, and the production of complete and partial flow datasets. Section IV presents the performance comparison of the models when applied to complete and partial flows, focusing on precision, recall, and F1-score metrics. Section V interprets the results, explores the study’s limitations, discusses prior studies that have addressed similar challenges. The paper concludes in Section VI.
II Background
II-A Readiness Criteria for Flow Categorization
In this paper, we opt to investigate the performance of machine learning algorithms (MLA) where the flows are considered ready for categorisation based on two separate thresholds:
  • First we investigate the performance of MLAs under a packet count threshold, focusing on the quantitative aspects of network communication.
  • Subsequently, we explore the efficacy of these algorithms under a time window threshold, emphasizing the temporal dynamics of network flows.
This dual-pronged approach allows us to comprehensively assess the behavior and effectiveness of MLAs in scenarios that closely mirror real-world conditions.
By separately analyzing these two threshold mechanisms, we aim to provide a more detailed and nuanced understanding of MLAs in real-time anomaly detection. This methodology aligns with our objective to improve the applicability and accuracy of anomaly detection systems, ensuring they are well-equipped for the challenges presented by contemporary network environments.
II-B Dataset
In this research, we leverage the CICIDS-2017 dataset [CICIDS2017], an extensive collection of labeled network traffic flows. This dataset, gathered from July 3 to July 7, 2017, is detailed across multiple PCAP files, segmented by each day of the week from Monday to Friday. Specifically, Monday’s dataset exclusively features benign traffic, while the data from Tuesday to Friday encompasses a varied mix of both benign and malicious traffic patterns, including DoS/DDoS, Port Scan, Brute Force, and Infiltration events, with each day presenting distinct types of attacks.
Due to the complexity of the dataset’s structure, producing a day-wise evaluation that is both comprehensive and accessible presents considerable challenges. Therefore, our analysis is specifically concentrated on the segment of network traffic from Wednesday, notable for its wide range of Denial-of-Service (DoS) attacks. This focus allows for a detailed examination of typical DoS attacks—such as DoS Hulk, DoS GoldenEye, DoS Slowloris, and DoS Slowhttptest—alongside the Heartbleed attack.
[…]
II-C ML Algorithm
In this study, we investigate the effects of incomplete flow information on network anomaly detection by employing the Random Forest algorithm, a technique widely recognized and frequently utilized in machine learning. RF, an ensemble learning method, builds numerous decision trees during the training phase and aggregates their predictions (in classification, this means taking the mode of the classes predicted by individual trees). This process not only boosts overall predictive accuracy but also helps prevent overfitting.
Opting for the RF algorithm enables us to draw on a method validated in related research [CICIDS2017, engelen2021, lanvin2023], thereby allowing us to uncover the broader implications of utilizing complete versus partial flow information. This approach facilitates the critical evaluation of the relevance and applicability of existing results within the context of real-time network security analytics.
III Methodology
III-A Raw Data Preprocessing
[…]
III-B Flow Metering
For the critical task of processing the PCAP file, we utilized NFStream [AOUINI2022108719], a Python-based tool designed for efficient, flexible, and detailed data processing in network analysis. NFStream is particularly effective at converting raw network traffic traces into structured data suitable for advanced analytics. It features robust flow measurement and feature computation capabilities, underpinned by a versatile architecture. This adaptability is notably evident in its NFPlugin component [AOUINI2022108719], which allows for the integration of custom network functionalities. In our research, this adaptability was crucial, enabling the incorporation of a precise flow labeling methodology directly into the analytical framework.
Our labeling mechanism is designed to classify network flows into distinct categories based on predefined criteria that reflect the attack patterns and benign behaviors documented by the original dataset contributors [CICIDS2017]. It systematically assigns labels by evaluating each flow’s metadata, including source and destination IP addresses, port numbers, protocol types, packet payload sizes, and temporal initiation details. Additionally, this mechanism includes a feature that, upon activation, reverses the flow’s direction by swapping the source and destination parameters. This ensures accurate flow labeling, particularly in scenarios where packets from the same flow may be segmented into subflows.
III-C Preliminary Measurement
The objective of our preliminary measurement was to closely emulate the conditions of the original CICIDS-2017 dataset, aiding in the calibration of flow measurement settings for complete flow records. Accordingly, we adjusted the idle and active timeout settings in NFStream to 60 and 120 seconds, respectively, a change from the default parameters of 120 and 1800 seconds.
From the results obtained, several conclusions can be drawn:
  • We observed a considerable number of flows with zero packet payload (ZPL), a characteristic not associated with any of the attacks identified in the Wednesday traffic trace, as detailed in Table I.
  • A considerable number of flows featured an unusually high count of FIN and RST packets. Table II summarizes these counts, specifically highlighting flows with more than two FIN or RST flags. Interestingly, the attacks observed on Wednesday (various DoS attacks and Heartbleed) are not characterized by an increased number of packets with TCP FIN or RST flags. This finding suggests that the packet traces may capture not only the attack signatures but also the repercussions, such as servers potentially saturated and beginning to terminate connections by sending packets with these TCP flags.
  • The Heartbleed attack appears as one prolonged attack that is segmented into multiple flow records by the active timeout setting.
  • A pattern of repeated flows was identified across the dataset, distinguished using a five-tuple of source and destination IP and port numbers, alongside the protocol identifier.
  • Numerous flows exhibited a Packet Inter-Arrival Time (PIAT) marginally below the idle timeout, set at 60 seconds. These instances might suggest that separate flows were amalgamated into a single flow due to the idle timeout configuration. Alternatively, this behavior could imply that some attacks were deliberately kept active, with packets sent just before the expiration of the timeout.
Complementary insights supporting the observations detailed above are available as a digital artifact, accessible via [github-repo].
TABLE I: Anomaly Breakdown
Flow Type Count ZPL Count
DoS GoldenEye 7 916 870
DoS Hulk 158 027 594
DoS Slowhttptest 3 010 3 088
DoS Slowloris 5 192 1 805
Heartbleed 11
$\sum$ Attacks 174 156 6 357
BENIGN 310 381 13 580
$\sum$ Total 484 537 19 937
TABLE II: Wednesday Traffic TCP FIN/RST Breakdown
Metric Value
BENIGN FIN $>$ 2 3 008
Attack FIN $>$ 2 5 819
Total FIN $>$ 2 8 827
BENIGN RST $>$ 2 2 273
Attack RST $>$ 2 38 559
Total RST $>$ 2 40 832
III-D Producing Complete Flows
Following the insights from Section III-C, we configured NFStream to generate complete flows with these settings:
  • The initial timeout is maintained at 60 seconds, falling within the range of typical settings for Linux Kernel netfilter and IPv4 TCP-specific networking configurations.
  • The active timeout is set to 18,000 seconds (5 hours) to prevent the segmentation of exceptionally long flows due to timeout expiration.
  • A TCP FIN/RST flag-based flow expiration policy has been implemented, which terminates flows at the first detection of either a FIN or RST flag, whether at flow initiation or upon update. This approach aims to exclude the aftermath of attacks that manifest as connection terminations. As a result, residual flow fragments, typically single-packet flows marked by a FIN or RST flag, are excluded.
  • The flow start time is now incorporated into the flow ID hash, enhancing the ability to match complete flows with their corresponding partial counterparts. Potential duplicate flow hash entries (despite this addition to the six-tuple used for unique flow identification) are discarded.
  • All flows with zero packet payloads (ZPL) are excluded.
  • The Heartbleed attack is omitted due to an insufficient number of samples for meaningful classification performance evaluation.
The distribution of the refined dataset, containing complete flows generated in accordance with the aforementioned methodology, is detailed in Table III. Complementary insights supporting the decisions detailed above are available as a digital artifact, accessible via [github-repo].
TABLE III: Complete Flow Dataset Breakdown
Flow Type Count
DoS GoldenEye 7,917
DoS Hulk 158,680
DoS Slowhttptest 3,707
DoS Slowloris 5,683
$\sum$ Attacks 175,987
BENIGN 326,363
$\sum$ Total 502,350
Compared to the data in Table I, Table III presents a marginally higher number of flow records for specific types. This increase in Table III, even after excluding flows with ZPL, results from our TCP flow expiration policy, which segments flows with multiple, irregular sequences of TCP FIN and RST flags into separate subflows. To mitigate any potential bias arising from such segmentation, post-processing steps are undertaken. In this phase, partial flows are matched with their complete counterparts using their six-tuple identification. This ensures that the initial segments of partial flows are accurately compared and aligned with the corresponding complete flows.
III-E Producing Partial Flows
In our partial flow measurement approach, we implemented two key mechanisms:
III-E1 Packet count
the first mechanism targets the precision of packet counts within flows: it ensures that only flows with an exact number of packets, as predefined, are retained for analysis. This strict selection criterion allows us to focus sharply on specific data exchange patterns, eliminating any flow that does not meet the exact packet count threshold. The range for this measurement was selected as $N_{pc}=\{2,3,4,\ldots,20\}$.
III-E2 Flow duration
the second mechanism deals with the duration of flows. Unlike the strict packet count approach, this mechanism allows for a degree of flexibility by retaining flows whose durations fall within a $\pm$20range of a specified target duration. This variance accommodates the dynamic nature of network traffic, ensuring that our analysis remains robust without being excessively restrictive. By applying this range, we acknowledge the natural fluctuations in flow durations while still maintaining a focus on our targeted temporal scope. The range for this measurement was selected as $N_{fd}=\{5,10,50,100,150,300,500,1000,5000,10000,15000,20000\}$ milliseconds.
In the process of producing partial flows, direct flow labeling has not been conducted. Instead, our focus was on identifying the complete flow counterparts for each partial flow using six-tuple hashes, subsequently assigning the corresponding labels from the complete flows to their partial counterparts. During this process, we also excluded any partial flows that did not meet the minimum value criteria specified in $N_{pc}$ and $N_{fd}$ for each respective dataset.
These methodologies, by specifically targeting packet count and flow duration, facilitate a nuanced analysis of the impact of partial flows on anomaly detection. This approach provides a comparative insight into how partial flow information influences the detection process relative to analyses based on complete flow information.
IV Results
IV-A Packet Count-based Evaluation
TABLE IV: Distribution of Complete and Partial Network Flows by Packet Count Thresholds
DS TOTAL BENIGN ANOMALY Anomaly Type Count Min Dur. [ms] Mean Dur. [ms] Max Dur. [ms]
CF 502 350 326 363 175 987 DoS GoldenEye 7 917 0 11 028.96 106 793
DoS Hulk 158 680 0 693.68 128 843
DoS Slowhttptest 3 707 0 9 562.98 167 903
DoS Slowloris 5 683 0 33 548.82 105 745
PC=2 500 493 324 508 175 985 DoS GoldenEye 7 917 0 130.35 58 333
DoS Hulk 158 680 0 456.34 32 096
DoS Slowhttptest 3 705 0 152.66 36 864
DoS Slowloris 5 683 0 168.72 1 000
PC=3 290 145 119 229 170 916 DoS GoldenEye 7 568 0 204.30 58 333
DoS Hulk 158 559 0 528.67 64 257
DoS Slowhttptest 2 419 0 3 055.00 43 264
DoS Slowloris 2 370 0 3 403.97 57 769
PC=4 271 751 101 923 169 828 DoS GoldenEye 7 567 0 233.50 48 079
DoS Hulk 158 557 0 529.51 65 721
DoS Slowhttptest 1 469 0 1 415.11 98 560
DoS Slowloris 2 235 0 908.58 3 004
PC=5 263 319 93 547 169 772 DoS GoldenEye 7 567 0 255.96 48 080
DoS Hulk 158 557 0 552.16 65 721
DoS Slowhttptest 1 413 1 2 068.20 15 036
DoS Slowloris 2 235 0 989.58 3 005
PC=6 261 477 92 476 169 001 DoS GoldenEye 7 567 0 4 773.53 52 865
DoS Hulk 158 557 0 637.06 66 719
DoS Slowhttptest 679 8 6 174.75 31 079
DoS Slowloris 2 198 1 2 466.67 53 080
PC=7 260 091 91 375 168 716 DoS GoldenEye 7 567 0 5 415.78 63 095
DoS Hulk 158 461 0 659.39 66 719
DoS Slowhttptest 631 9 8 649.36 63 160
DoS Slowloris 2 057 2 3 017.98 54 078
PC=8 251 718 87 891 163 827 DoS GoldenEye 7 567 1 5 743.47 63 095
DoS Hulk 153 941 0 688.26 68 723
DoS Slowhttptest 458 11 11 213.53 63 160
DoS Slowloris 1 861 3 2 364.51 54 079
PC=9 222 951 85 492 137 459 DoS GoldenEye 7 411 1 6 757.70 63 095
DoS Hulk 127 790 0 784.01 68 723
DoS Slowhttptest 415 17 12 643.28 63 161
DoS Slowloris 1 843 3 3 583.45 59 080
PC=10 141 273 84 555 56 718 DoS GoldenEye 6 183 1 6 603.48 63 096
DoS Hulk 48 451 1 1 602.42 72 731
DoS Slowhttptest 246 2 603 20 056.63 63 161
DoS Slowloris 1 838 21 5 688.29 59 080
PC=11 102 837 82 489 20 348 DoS GoldenEye 5 585 2 6 550.69 63 096
DoS Hulk 12 706 2 2 787.30 72 731
DoS Slowhttptest 239 2 655 22 290.57 63 364
DoS Slowloris 1 818 22 9 709.68 50 039
PC=12 91 649 80 104 11 545 DoS GoldenEye 5 253 2 7 293.73 68 586
DoS Hulk 4 245 10 2 991.83 80 753
DoS Slowhttptest 236 7 619 25 932.32 63 568
DoS Slowloris 1 811 22 18 236.57 25 680
PC=13 83 175 75 062 8 113 DoS GoldenEye 4 366 2 8 688.62 68 586
DoS Hulk 1 716 10 4 343.46 80 754
DoS Slowhttptest 234 13 088 29 969.69 63 976
DoS Slowloris 1 797 22 25 719.71 31 622
PC=14 75 628 70 410 5 218 DoS GoldenEye 2 560 3 9 709.52 68 586
DoS Hulk 642 161 8 026.42 96 779
DoS Slowhttptest 234 13 088 33 832.93 64 792
DoS Slowloris 1 782 212 30 787.90 32 624
PC=15 69 081 65 645 3 436 DoS GoldenEye 1 246 5 11 117.28 68 587
DoS Hulk 181 161 18 406.82 96 779
DoS Slowhttptest 234 13 089 41 071.45 66 428
DoS Slowloris 1 775 222 42 131.41 51 365
PC=16 64 929 62 137 2 792 DoS GoldenEye 743 10 13 591.32 106 793
DoS Hulk 61 1 612 57 281.75 128 842
DoS Slowhttptest 234 13 110 55 096.79 107 103
DoS Slowloris 1 754 222 70 441.35 102 735
PC=17 61 237 58 779 2 458 DoS GoldenEye 487 12 16 609.48 106 793
DoS Hulk 43 1 612 77 516.74 128 843
DoS Slowhttptest 234 13 110 67 436.24 107 773
DoS Slowloris 1 694 235 87 924.51 103 735
PC=18 57 217 55 799 1 418 DoS GoldenEye 313 1 661 18 334.05 105 645
DoS Hulk 13 1 650 14 255.54 15 483
DoS Slowhttptest 202 13 150 74 849.35 107 773
DoS Slowloris 890 272 104 104.38 105 745
PC=19 54 154 53 267 887 DoS GoldenEye 206 2 746 21 941.75 105 645
DoS Hulk 13 2 051 14 449.23 16 355
DoS Slowhttptest 174 18 112 85 884.79 111 750
DoS Slowloris 494 1 675 105 045.52 105 745
PC=20 51 267 50 961 306 DoS GoldenEye 150 2 746 24 907.24 73 588
DoS Hulk 13 2 055 24 756.85 27 741
DoS Slowhttptest 141 52 390 100 305.45 119 800
DoS Slowloris 2 2 250 6 966.50 11 683
(a) Binary classification performance across packet count scenarios.
(b) Multi-class classification performance across packet count scenarios.
Figure 1: Comparative analysis of binary and multi-class classification performance across different packet count scenarios.
IV-A1 CF and PF Dataset Distribution
Table IV presents the dataset distributions for packet count-based evaluation. The table starts with the ”Complete Flows” (CF) category, which encompasses the ground truth dataset (see Section III-D) and then breaks down into partial flows labeled as PC=N, where N ranges from 2 to 20, representing datasets characterized by specific packet counts. For each category, Table IV lists the total number of flows, classified into benign and anomalous types, with anomalies further subdivided into distinct types of DoS attacks. The table also includes metrics on the minimum, mean, and maximum durations (in milliseconds) for these categories.
Analysis of Table IV reveals a significant decrease in the total number of flows as the packet count per flow increases. The dataset starts with 502,350 network flows in the CF category, predominantly benign (326,363), with 175,987 flows categorized as anomalous, spanning four types of DoS attacks. In contrast, at PC=20, the dataset contains only 51,267 flows, with 50,961 being benign and merely 306 classified as anomalous.
This trend continues across specific attack types: as packet counts in flows increase, the number of records correspondingly decreases, suggesting that many attacks transmit only a few packets. Moreover, there is an observable increase in the maximum duration of attacks as packet count per flow increases, especially notable in Slowhttptest and Slowloris attacks, which are known for prolonged durations. Thus, the packet count per flow crucially influences the observed characteristics of network attacks, with higher packet counts typically associated with longer durations. Conversely, many attacks consist of flows with few packets, indicating that extracted flow statistics might lack sufficient information for precise anomaly detection, complicating effective detection mechanisms.
IV-A2 CF vs. PF Performance Comparison
Figure 1 illustrates the precision, F1-score, and recall for both binary (Figure 1a) and multi-class (Figure 1b) classification under three scenarios:
  • CF used for both training and testing,
  • PC=N for both training and testing, and
  • CF used for training with PC=N for testing.
Note that PC ranges from 2 to 17, unlike in Table IV, because PCs 18, 19, and 20 lacked enough samples for meaningful comparison. Additionally, before training, we selected only those flows from the CF dataset that intersect with the flow hash used as identification in the respective partial flow dataset.
From Figure 1a, it is observed that both CF train/CF test and N train/N test scenarios achieve sufficient performance starting at PC=2, though performance is slightly lower at minimal PC values in the N train/N test scenario. However, the CF train/N test scenario highlights the negative impacts of insufficient information due to lower packet counts, indicating that at least 8 packets are necessary in the N test to achieve usable performance using CF train model, with some minor fluctuations at higher packet counts.
The multi-class classification in Figure 1b similarly shows that low packet counts negatively affect the N train/N test scenario, requiring at least 6 packets to achieve acceptable performance. Additionally, even the highest packet count seems inadequate for stable performance across precision, F1-score, and recall in the CF train/N test scenario, underscoring the challenges of anomaly detection with limited data when only completed flows have been used for training.
IV-B Flow Duration-based Evaluation
TABLE V: Distribution of Complete and Partial Network Flows by Flow Duration Thresholds
DS TOTAL BENIGN ANOMALY Anomaly Type Count Min Pckts. Mean Pckts. Max Pckts.
CF 502 350 326 363 175 987 DoS GoldenEye 7 917 2 12.14 30
DoS Hulk 158 680 2 9.20 27
DoS Slowhttptest 3 707 1 4.72 35
DoS Slowloris 5 683 2 7.36 27
FD=5ms 7 693 729 6 964 DoS GoldenEye 270 3 7.11 15
DoS Hulk 5 917 3 4.03 10
DoS Slowhttptest 777 2 3.98 4
FD=10ms 82 521 1 093 81 428 DoS GoldenEye 107 3 9.22 14
DoS Hulk 81 321 2 3.16 11
FD=50ms 53 984 48 309 5 675 DoS Hulk 5 675 3 7.37 11
FD=100ms 46 776 42 783 3 993 DoS Hulk 3 993 4 7.27 11
FD=150ms 104 191 37 334 66 857 DoS Hulk 66 738 6 7.81 11
DoS Slowhttptest 119 2 2.44 4
FD=300ms 31 784 25 597 6 187 DoS Hulk 6 187 5 7.16 12
FD=500ms 20 237 17 266 2 971 DoS Hulk 2 832 8 8.18 13
DoS Slowloris 139 5 6.02 7
FD=1000ms 24 545 9 606 14 939 DoS GoldenEye 442 2 4.08 8
DoS Hulk 13 856 2 5.56 14
DoS Slowhttptest 152 2 3.16 8
DoS Slowloris 489 2 4.22 7
FD=5000ms 15 953 13 714 2 239 DoS GoldenEye 2 239 6 8.37 18
FD=10000ms 37 060 31 689 5 371 DoS GoldenEye 5 371 6 12.10 24
FD=15000ms 8 763 4 243 4 520 DoS GoldenEye 510 2 13.11 27
DoS Hulk 3 314 2 2.06 19
DoS Slowhttptest 138 5 6.43 13
DoS Slowloris 558 6 12.68 25
FD=20000ms 30 545 29 864 681 DoS GoldenEye 341 8 14.11 24
DoS Slowhttptest 340 3 5.01 19
(a) Binary classification performance across flow duration scenarios.
(b) Multi-class classification performance across flow duration scenarios.
Figure 2: Comparative analysis of binary and multi-class classification performance across different flow duration scenarios.
IV-B1 CF and PF Dataset Distribution
Similar to Table IV, Table V outlines the dataset distributions for flow duration-based evaluation. This methodical categorization delineates a transition from ”Complete Flows” (CF) to progressively more granular ”Partial Flows” (PF), identified as FD=N, where N encompasses a range of specified durations ${5,10,50,100,150,300,500,1000,5000,10000,15000,20000}$ milliseconds. Each category within Table V specifies the total number of flows, distinguishing between benign and anomalous flows, with the latter further divided into distinct DoS attack types. Contrasting with Table IV, which detailed flow durations, this table provides insights on the minimum, mean, and maximum packet counts for these categories.
Upon examining Table V, a similar trend emerges as was seen with packet count-based datasets; however, when flow duration thresholds define partial flow generation, there’s a dramatic drop in flow counts for each dataset. This decrease persists even though our methodology accommodates flows with durations within $\pm$20of the threshold values, as detailed in Section III-E.
The statistics in flow count with various flow durations underscores the challenge of capturing meaningful data within constricted time frames, particularly for anomaly detection where comprehensive flow information is crucial. Across all flow we see a fluctuating count of flows per most cateogries. For example, for DoS Hulk with FD=10ms, there are 81,321 flows, which decreases to 66,738 flows at FD=150ms, drops further to 2,832 at FD=500ms, and then unexpectedly rises again to 13,856 at FD=1000ms. Contrary to expectations, the flow count does not monotonically increase with longer durations.
The irregular trend could result from the inherent nature of network traffic during a DoS Hulk attack, where periods of intense activity are interspersed with lulls. As the duration threshold increases, the dataset may initially miss capturing these shorter bursts of activity, resulting in a lower count. Yet as the threshold extends further, it may again encompass subsequent waves of attack traffic, thus accounting for the rise in flow count.
IV-B2 CF vs. PF Performance Comparison
Figure 2 depicts the performance metrics of precision, F1-score, and recall for binary and multi-class classifications across a range of FD thresholds. For the sake of this analysis, only FD=5, 10, 150, 500, 1000, 15000, and 20000 ms are included, as other durations did not have enough data to yield statistically significant results. Analogous to the packet count-based evaluation, prior to training, we exclusively selected flows from the CF dataset that correspond with the flow hash identifiers present within the respective partial flow dataset.
Figure 2a confirms the trend observed in packet count evaluations. When the training and testing datasets match (CF train/CF test and N train/N test), binary classification performance is strong and stable. However, the model’s ability to accurately classify benign and anomalous flows is compromised in the CF train/N test scenario, reflecting the challenge of applying a model trained on complete flow data to partial flows characterized by various durations.
In Figure 2b, this issue is further exacerbated in the multi-class classification scenario. Here, even the N train/N test performance demonstrates a notable decline, particularly at FD=1000ms, suggesting that as with certain flow durations set as threshold, the model’s ability to distinguish between multiple classes of traffic—beyond benign and anomalous—becomes is unreliable. This decline in performance may be attributed to the reduced information content in flows with various duration, which does not adequately capture the complexities of multi-class traffic patterns.
The performance dip in the CF train/N test scenario for multi-class classification is even more pronounced, reinforcing the notion that comprehensive flow data is crucial for developing robust models. The precision, F1-score, and recall all experience a steep drop when the model trained on complete data is tested on partial data. This emphasizes the model’s dependence on the quantity and quality of training data and the challenge of applying it to substantially different test conditions.
V Discussion
This study expands our packet count- and flow duration-based evaluations with additional digital artifacts [github-repo]. These materials provide a more comprehensive set of metrics, including accuracy, balanced accuracy, and confusion matrices. Moreover, they offer visualizations that extend beyond the scope of those included in the main body of this paper.
V-A Implications
Our findings confirm that models trained and tested on the same data distribution (CF or N) exhibit commendable performance. However, a decrement in packet count introduces a performance dip, suggesting that a model with less data may struggle to generalize well. This effect is particularly pronounced in cases where models trained on complete data (CF) are applied to partial data (PC=N), highlighting potential overfitting to the complete data’s characteristics.
The analysis indicates that while higher packet counts are associated with extended flow durations, attacks with fewer packets could present challenges for precise anomaly detection due to insufficient detail in flow statistics. For effective binary classification, a threshold of at least 7 packets in the test data is essential to ensure performance reliability when classifying using CF dataset (CF train/N test), while for multi-class classification, at least 6 packets are necessary when trained and tested on the same dataset (N train/N test).
Fluctuations in performance across various packet counts emphasize the significance of selecting an optimal packet count threshold for maintaining classification accuracy. When training and testing datasets are aligned, performance remains stable. Yet, the application of models developed on complete flows to partial flows poses a significant challenge, underscoring the importance of data quality and volume in training robust anomaly detection models.
Additionally, the non-linear trends observed across different flow duration thresholds highlight the complexity inherent in network attack profiling, which cannot be solely based on duration metrics. The adaptive and dynamic nature of such attacks necessitates a more nuanced detection approach that can respond to varied attack patterns and evasion techniques.
Interestingly, while a specific packet count threshold can be identified that yields stable classification performance, no such threshold emerged for flow duration. This underscores the challenge in defining a one-size-fits-all threshold for flow duration-based classification.
Overall, the research presented here sheds light on the complex interplay between data distribution and classification efficacy, leading to essential considerations for deploying network security measures. Future models must be designed with the flexibility to adjust to varying levels of traffic information, ensuring that anomaly detection remains both robust and reliable.
V-B Limitations
While this research provides valuable insights into anomaly detection using Random Forest models, it is not without limitations that may affect the generalizability and application of the findings. First, the study does not extend to the examination of flows with variable lengths. Our analysis primarily focuses on flows characterized by a fixed number of packets, assessing them from both packet count and flow duration perspectives. This approach limits our ability to fully understand the dynamics of more complex, variable-length network flows that often occur in real-world environments. The fixed-length assessment may overlook nuances that appear only in longer or dynamically changing sequences, potentially affecting the model’s ability to generalize across different network conditions.
Secondly, the current study focuses exclusively on the first $N$ packets per flow. This methodology may miss crucial information contained in subsequent packets of a flow, especially in longer or continuous data streams. It might be beneficial to assess the flows through a sliding window approach or a sequence of sub-flows, which could capture changes and anomalies that develop beyond the initial packets.
Additionally, analyzing partial flows with packet counts starting from 2 may not adequately address the complexities of modern protocols such as TLS and QUIC, where connection or session setup alone may require more packets than the lower range of our study. Focusing on such an early packet range might overlook crucial information for classification that appears only during or after the completion of these setups.
Packet count thresholds also face challenges with hardware offloading technologies, which, after a certain number of packets, can shift the maintenance of flows to fast path processing. Such hardware-level interventions can obscure or completely hide parts of the flow from observation, complicating detection strategies that rely on early-stage categorization.
Furthemore, our method of generating partial flows based on flow duration with different thresholds highlights a challenge. While allowing a degree of flexibility by retaining flows whose durations fall within a ±$\pm 20\%$ range of a specified target duration, a dramatic drop in flow counts for each dataset indicates that our approach may still be too restrictive. A more refined methodology is needed to better assess and accommodate the variable nature of flow durations, ensuring that our anomaly detection framework can effectively handle the diverse and dynamic characteristics of network traffic.
Lastly, this study examines the performance of the RF algorithm. To assess generalizability, we also evaluated Decision Trees (DT), whose codes are accessible via [github-repo]. While DT approached the performance of RF, generally, RF outperformed DT in our evaluations. However, an extended analysis involving a broader range of various MLAs is needed to draw more targeted conclusions about the effectiveness of different models.
V-C Related Work
[…]
While these studies contribute significantly to the field, they often do not comprehensively examine crucial aspects such as packet count and flow duration, both essential for ensuring robust and reliable anomaly detection in real-world scenarios. Our research contrasts the use of packet count and flow duration thresholds for early-stage anomaly detection, establishing a foundational understanding of the trade-offs between various threshold strategies for network traffic analysis. This nuanced approach allows our study to provide practical insights that can significantly influence real-time network security measures.
V-D Comparison with Relevant Research
[…]
VI Conclusion
This study examined the efficacy of the Random Forest model in the context of anomaly detection, applying it to both complete and partial network flow data derived from a refined version of the CICIDS2017 dataset. Our findings underscore the necessity of using at least 7 packets to achieve reliable anomaly detection performance, contrasting with recent studies that suggest fewer packets may be sufficient.
Our research contributes to the development of machine learning models that accurately reflect the challenges of analyzing partial flow data in real-time network environments. By aligning model development and validation with the operational dynamics inherent in real-time anomaly detection, this study helps bridge critical gaps in current research methodologies.
Future work will focus on refining data preprocessing techniques to handle variable-length data more effectively and exploring the integration of more complex machine learning models that can automatically adjust to the dynamic nature of network traffic, potentially improving the accuracy and responsiveness of anomaly detection systems.
Acknowledgement
This work was supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. Supported by the ÚNKP-23-5-BME-461 New National Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. The work presented in this paper was supported by project no. TKP2021-NVA-02. Project no. TKP2021-NVA-02 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.
[…]